OpenStudy (anonymous):

solve for x... please i'm in a middle of a test 13^(-10x)=8^(x-5)

OpenStudy (anonymous):

awesome man, thank you so much..

OpenStudy (anonymous):

Isn't it forbidden to give answers for a test?

OpenStudy (campbell_st):

I don't know... but where is the answer...

OpenStudy (anonymous):

Good job. You got rid of it.

OpenStudy (campbell_st):

lol... do you need it back..?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I am glad you got rid of it.

OpenStudy (campbell_st):

why is that...?

OpenStudy (campbell_st):

we cheat all the time in my grade 8 class... the teacher will tell us some answers...

OpenStudy (anonymous):

Hmm... I don't agree with cheating but okay, put it back up if you want...

OpenStudy (campbell_st):

well... all test tell you... is what you don't know... and he doesn't know log laws...

OpenStudy (anonymous):

hey, so i have this other question; 5^(x+7)=9^(-5x) would x= @campbell_st